Project

General

Profile

12/02/2013 Minutes

Attendance

Operations

Smooth running over the weekend. Beam was down for 4 hours yesterday. Beam pitch tones are normal so we are not Low Intensity running.
There was a timing issue over the weekend. TDU that we're using to timestamp our spills shifted by 52 clock ticks. That's 812ns.
Shifted around noon on Thursday and before midnight on Friday 4092 clock ticks (64us) until Saturday.
Unsure why this happened. TDU not in any error state. Issue new time syncs did not recover things. Only solution was scrubbing the unit and returning to normal state. It's not supposed to do this so we will want to correct for this in analysis data.
Tricia suggests we need a Time Calibrator to correct for this offline - depends on how often this occurs.
Saturday there was an email to Data Quality about this very problem.

Action Item: Evan will get precise timings for this occurrence. There is an ECL entry. An update on the precise times and shift is forthcoming. But analysis efforts need to correct for this. A "hack" is the temporary solution for the run this occurred during.

APD installations continue today. APDs are being installed on di-block 5 DCMs on the top. Di-block 4 top is missing one DCM.
Di-block 4 we are installing no-primer APDs. Baked, older APDs on di-block 5. Only installing on the top but possibly will be able to start installing on di-block 3 later this week

Processing

Action Item: Robert has been alerted that File Transfer has stopped due some unknown failure. We're not processing anymore.

OnMon

Michael would like to stop trying to process Nearline plots for NDOS - it's cumbersome and time consuming and not clear the effort is actually being used or is critical.